OpenStudy (anonymous):

STOP THIS NON-SENSE NOW !!!!!!!!!!! Graph the system of inequalities. Then use your graph to identify the point that represents a solution to the system. x > 3 y ≤ 2x – 5 A. (5, 6) B. (4, –1) C. (–3, –1) D. (1, 11)

OpenStudy (anonymous):

wut

OpenStudy (anonymous):

i copied the first part on accident i meant to just put da question

OpenStudy (anonymous):

@gurrilabill

OpenStudy (anonymous):

so can u help

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

actually idk sorry

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

do you really need to graph, or do you want an easier way to solve this problem ?

OpenStudy (anonymous):

easier way

OpenStudy (texaschic101):

try subbing in your points into the inequalities. Look at x > 3.....this inequality rules out answer choices C and D because x has to be greater then 3. y <= 2x - 5 lets check A (5,6) x = 5 and y = 6 now sub 6 <= 2(5) - 5 6 <= 10 - 5 6 <= 5 (incorrect)...its not A y <= 2x - 5 lets check B (4,-1) x = 4 and y = -1 now sub -1 <= 2(4) - 5 -1 <= 8 - 5 -1 <= 3 (correct) B is your answer because the points satisfy both inequalities.